OpenStudy (anonymous):

Use the given information about a polynomial whose coefficients are real numbers to find the remaining zeros of the polynomial. Degree: 6 zeros: -19+PIE[i], -9+7i, 20+6i^3

OpenStudy (anonymous):

\[-19-\pi i\] and also the conjuates of the other two

OpenStudy (anonymous):

This is the answer?

OpenStudy (anonymous):

you have to take the conjugate of the other two the conjugate of \[a+bi\] is \[a-bi\]

OpenStudy (anonymous):

also note that \[20+6e^3=20-6i\] because \[i^3=-i\]

OpenStudy (anonymous):

\[20+6i^3=20-6i\]